# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v5

@@ Abstract
This paper addresses the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Gossip algorithms are fully decentralized, relying only on local neighbor-to-neighbor communication, making them well-suited for situations where communication is constrained. A fundamental challenge in existing mean-based gossip algorithms is their vulnerability to malicious or corrupted nodes. In this paper, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. More specifically, we propose a novel gossip algorithm for rank estimation, referred to as GoRank, and leverage it to design a gossip procedure dedicated to trimmed mean estimation, coined GoTrim. In addition to a detailed description of the proposed methods, a key contribution of our work is a precise convergence analysis: we establish an $\mathcal{O}(1/t)$ rate for rank estimation and an $\mathcal{O}((\log t)/\smash{\sqrt{t}})$ rate for trimmed mean estimation, where by $t$ is meant the number of iterations. Moreover, we provide a breakdown point analysis of GoTrim. We empirically validate our theoretical results through experiments on diverse network topologies, data distributions and contamination schemes.
@@ 1 Introduction
$\bullet$ We introduce GoTrim, the first gossip algorithm for trimmed mean estimation which can be paired with any ranking algorithm. GoTrim does not rely on strong graph topology assumptions—a common restriction in robust gossip algorithms. We prove a competitive convergence rate of order $\mathcal{O}((\log t)/\smash{\sqrt{t}})$ and provide a breakdown point analysis, demonstrating the robustness of its estimate.
@@
Figure 1: Illustration of a GoTrim’s iteration paired with GoRank on a communication graph $\mathcal{G}=(V,E)$ Initially, each node $k\in V$ stores the observation $X_{k}$.
@@ 4 GoTrim – A Gossip Algorithm for Trimmed Means Estimation
In this section, we present GoTrim, a gossip algorithm for trimmed means estimation. We establish a convergence in $\mathcal{O}(\log(t)/\smash{\sqrt{t}})$ with constants depending on the network and data distribution.
@@ 4.1 Algorithm – Convergence Analysis
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{1}{\gamma_{k}\sqrt{ct}}\right)\cdot\sigma_{n}(r_{k}),$
@@
Further details and the complete proof are provided in Section D of the Supplementary Material. Lemma 1 establishes that, for each node $k$, the estimates of the weight $W_{k}$ converge in expectation to the true weight $w_{n,\alpha}(r_{k})$ at a rate of $\mathcal{O}\left(1/\gamma_{k}\smash{\sqrt{ct}}\right)$. In addition, this lemma suggests that points closer to the interval endpoints are subject to larger errors, as reflected in the constant $\gamma_{k}$, which is consistent with the intuition that these points require greater precision in estimating ranks.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm 2, and assume the ranking algorithm is Algorithm 1. Then, we have that for all $k\in[n]$, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>\log(tc^{-1})\right\}$,
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\mathcal{O}\left(\frac{\log(tc^{-1})}{\sqrt{c^{2}t}}\right)\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=(\sigma_{n}(r_{1})/\gamma_{1},\;\ldots,\;\sigma_{n}(r_{n})/\gamma_{n})$ and $c$, $\sigma_{n}(r_{k})$, $\gamma_{k}$ are the constants defined in Lemma 1.
See Section D of the Supplementary Material for the detailed proof. Theorem 3 shows that, for each node $k$, the estimate of the trimmed mean $Z_{k}$ converges in expectation to the true trimmed mean $\bar{x}_{\alpha}$ at a rate of $\mathcal{O}\left(\log(tc^{-1})/\smash{\sqrt{c^{2}t}}\right)$. While the presence of the $c^{2}$ term may appear pessimistic, empirical evidence suggests that the actual convergence may be faster in practice. The vector $\mathbf{K}$ acts as a rank-dependent mask over $\mathbf{X}$, modulating the contribution of each data point $X_{k}$ to the error bound. Specifically: (1) extreme values are more heavily penalized by the mask, as they come with better rank estimation; (2) values with ranks near the trimming interval endpoints are amplified, as small inaccuracies in rank estimation can lead to disproportionately larger errors.
@@ 5 Conclusion and Future Work
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and $\mathcal{O}((\log t)/\smash{\sqrt{t}})$, respectively, and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean. Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in the Supplementary Material; however, similar to Baseline++, its convergence analysis is left for future work. In addition, extending GoTrim for decentralized estimation of rank statistics [20], key tools in data analysis, e.g., for robust hypothesis testing, is another promising direction. See the Supplementary Material for further discussion.
@@ D.1 Proof of Lemma 1
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\sqrt{\frac{3}{ct}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}(1-2\alpha)},$ (20)
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|)\leq\sqrt{\frac{3}{ct}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}},$ (21)
The right-hand side of the inequality follows directly from Theorem 2 via an application of Markov’s inequality:
$\mathbb{P}[|R_{k}(t)-r_{k}|\geq\gamma_{k}]\leq\frac{\mathbb{E}[|R_{k}(t)-r_{k}|]}{\gamma_{k}}.$
@@ D.2 Proof of Theorem 3
Let us assume that $G$ is connected and non-bipartite. Then, for $\mathbf{Z}(t)$ defined GoTrim, and assuming the ranking algorithm is GoRank, we have that for all $k\in[n]$ :
@@
Moreover, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>\log(t/c)\right\}$, we have
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\left(\frac{3}{\sqrt{ct}}+\frac{\log(t/c)}{(ct-\log(t/c))^{1/2}}\right)\cdot{\frac{\sqrt{3}}{\sqrt{c}(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n}$ and $c$, $\sigma_{n}(r_{k})$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
$\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{C_{i}}{\sqrt{t}}\cdot|X_{i}|$ (28)
$\leq\frac{1}{\sqrt{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|,$ (29)
where $C_{i}=\sqrt{\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{i})}{\gamma_{i}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-\frac{\log(\sqrt{t/c})}{c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(\sqrt{t/c})\right\}$. We obtain for all $t>T^{*}$,
@@
Using Lemma 1, we derive for $s>1$, $\Delta{w}_{k}(s)\leq\frac{C_{k}}{\sqrt{s-1}}$. We obtain $\|\mathbf{S}(1)\|\leq\|\mathbf{C}\odot\mathbf{X}\|$ and for $s>1$,
$\|\mathbf{S}(s)\|\leq\sqrt{\sum_{k=1}^{n}\frac{C_{k}^{2}X_{k}^{2}}{(s-1)}}\leq\frac{1}{\sqrt{s-1}}\|\mathbf{C}\odot\mathbf{X}\|.$
Applying this bound yields:
$\|\mathbf{R}(t)\|$ $\leq\lambda_{2}(2)^{t-T}\sum_{s=1}^{T}\lambda_{2}(2)^{T-s}\|\mathbf{S}(s)\|+(t-T)\lambda_{2}(2)\|\mathbf{S}(s)\|$
$\leq e^{-c_{2}(t-T)}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+(t-T)\frac{\lambda_{2}(2)}{\sqrt{T}}\|\mathbf{C}\odot\mathbf{X}\|$
$\leq\sqrt{\frac{c}{t}}\frac{1}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(\sqrt{t/c})}{({c_{2}t-\log(\sqrt{t/c}))^{1/2}}}\lambda_{2}(2)\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{1}{\sqrt{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|+\frac{2}{\sqrt{ct}}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\frac{1}{2}\log(t/c)}{({\frac{c}{2}t-\frac{1}{2}\log(t/c)})^{1/2}}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c_{2}=\left(1-\lambda_{2}(2)\right)>0$ with $\lambda_{2}(2)$ the second largest eigenvalue of $\mathbf{W}_{2}$ and the matrix $\mathbf{C}$ is defined as $\mathbf{C}=\left[\sqrt{\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{k})}{\gamma_{k}(1-2\alpha)}\right]_{k=1}^{n}.$ This bound simplifies to
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{3}{\sqrt{ct}}\|\mathbf{C}\odot\mathbf{X}\|+\frac{\log(t/c)}{(ct-\log(t/c))^{1/2}}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c=\frac{\lambda_{2}}{|E|}=2c_{2}$. Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|={\frac{\sqrt{3}}{\sqrt{c}(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})}{\gamma_{k}}\right]_{k=1}^{n},$ which completes the proof. ∎

### v6

@@ Abstract
This paper addresses the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Gossip algorithms are fully decentralized, relying only on local neighbor-to-neighbor communication, making them well-suited for situations where communication is constrained. A fundamental challenge in existing mean-based gossip algorithms is their vulnerability to malicious or corrupted nodes. In this paper, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. More specifically, we propose a novel gossip algorithm for rank estimation, referred to as GoRank, and leverage it to design a gossip procedure dedicated to trimmed mean estimation, coined GoTrim. In addition to a detailed description of the proposed methods, a key contribution of our work is a precise convergence analysis: we establish an $\mathcal{O}(1/t)$ rate for rank estimation and an $\mathcal{O}(1/{t})$ rate for trimmed mean estimation, where by $t$ is meant the number of iterations. Moreover, we provide a breakdown point analysis of GoTrim. We empirically validate our theoretical results through experiments on diverse network topologies, data distributions and contamination schemes.
@@ 1 Introduction
$\bullet$ We introduce GoTrim, the first gossip algorithm for trimmed mean estimation which can be paired with any ranking algorithm. GoTrim does not rely on strong graph topology assumptions—a common restriction in robust gossip algorithms. We prove a competitive convergence rate of order $\mathcal{O}(1/{t})$ and provide a breakdown point analysis, demonstrating the robustness of its estimate.
@@
Figure 1: Illustration of a GoTrim’s iteration paired with GoRank on a communication graph $\mathcal{G}=(V,E)$. Initially, each node $k\in V$ stores the observation $X_{k}$.
@@ 4 GoTrim – A Gossip Algorithm for Trimmed Means Estimation
In this section, we present GoTrim, a gossip algorithm for trimmed means estimation. We establish a convergence in $\mathcal{O}(1/{t})$ with constants depending on the network and data distribution.
@@ 4.1 Algorithm – Convergence Analysis
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\mathcal{O}\left(\frac{1}{\gamma_{k}^{2}{ct}}\right)\cdot\sigma_{n}(r_{k})^{2},$
@@
Further details and the complete proof are provided in Section D of the Supplementary Material. Lemma 1 establishes that, for each node $k$, the estimates of the weight $W_{k}$ converge in expectation to the true weight $w_{n,\alpha}(r_{k})$ at a rate of $\mathcal{O}\left(1/\gamma_{k}^{2}{ct}\right)$. In addition, this lemma suggests that points closer to the interval endpoints are subject to larger errors, as reflected in the constant $\gamma_{k}$, which is consistent with the intuition that these points require greater precision in estimating ranks.
@@
Let $\mathbf{Z}(t)$ be defined as in Algorithm 2, and assume the ranking algorithm is Algorithm 1. Then, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$,
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\mathcal{O}\left(\frac{1}{{c^{2}t}}\right)\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=(\sigma_{n}(r_{1})^{2}/\gamma_{1}^{2},\;\ldots,\;\sigma_{n}(r_{n})^{2}/\gamma_{n}^{2})$ and $c$, $\sigma_{n}(r_{k})$, $\gamma_{k}$ are the constants defined in Lemma 1.
See Section D of the Supplementary Material for the detailed proof. Theorem 3 shows that, for each node $k$, the estimate of the trimmed mean $Z_{k}$ converges in expectation to the true trimmed mean $\bar{x}_{\alpha}$ at a rate of $\mathcal{O}\left(1/{c^{2}t}\right)$. While the presence of the $c^{2}$ term may appear pessimistic, empirical evidence suggests that the actual convergence may be faster in practice. The vector $\mathbf{K}$ acts as a rank-dependent mask over $\mathbf{X}$, modulating the contribution of each data point $X_{k}$ to the error bound. Specifically: (1) extreme values are more heavily penalized by the mask, as they come with better rank estimation; (2) values with ranks near the trimming interval endpoints are amplified, as small inaccuracies in rank estimation can lead to disproportionately larger errors.
@@ 5 Conclusion and Future Work
We introduced and analyzed two novel gossip algorithms: GoRank for rank estimation and GoTrim for trimmed mean estimation. We proved convergence rates of $\mathcal{O}(1/t)$ and established robustness guarantees for GoTrim through breakdown point analysis. Empirical results show both methods perform well on large, poorly connected networks: GoRank quickly estimates ranks, and GoTrim is robust to outliers and improves on the naive mean. Although our analysis focuses on the synchronous setting, real-world systems are often asynchronous. We present the asynchronous version of GoRank in the Supplementary Material; however, similar to Baseline++, its convergence analysis is left for future work. In addition, extending GoTrim for decentralized estimation of rank statistics [20], key tools in data analysis, e.g., for robust hypothesis testing, is another promising direction. See the Supplementary Material for further discussion.
@@ D.1 Proof of Lemma 1
$\left|\mathbb{E}[W_{k}(t)]-w_{n,\alpha}(r_{k})\right|\leq\frac{3}{ct}\cdot\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}(1-2\alpha)},$ (20)
@@
$\left|\mathbb{P}(R_{k}(t)\in I_{n,\alpha})-\mathbb{I}_{\{r_{k}\in I_{n,\alpha}\}}\right|\leq\mathbb{P}(|R_{k}(t)-r_{k}|\geq\gamma_{k})\leq\frac{3}{ct}\cdot\frac{\sigma_{n}\left(r_{k}\right)^{2}}{\gamma_{k}^{2}},$ (21)
The right-hand side of the inequality follows directly from Lemma C.1 via an application of Markov’s inequality:
$\mathbb{P}[|R_{k}(t)-r_{k}|\geq\gamma_{k}]=\mathbb{P}[|R_{k}(t)-r_{k}|^{2}\geq\gamma_{k}^{2}]\leq\frac{\mathbb{E}[|R_{k}(t)-r_{k}|^{2}]}{\gamma_{k}^{2}}.$
@@ D.2 Proof of Theorem 3
Let $\mathbf{Z}(t)$ be defined GoTrim, and assuming the ranking algorithm is GoRank, we have that for all $k\in[n]$ :
@@
Moreover, for any $t>T^{*}=\min\left\{t>1\,\middle|\,ct>2\log(t)\right\}$, we have
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\left(\frac{5}{{ct}}+\frac{4}{ct-2\log(t)}\right)\cdot{\frac{{3}}{c(1-2\alpha)}}\cdot\|\mathbf{K}\odot\mathbf{X}\|,$
where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n}$ and $c$, $\sigma_{n}(r_{k})$ and $\gamma_{k}$ are constants defined in Lemma 1.
@@
$\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{C_{i}}{{t}}\cdot|X_{i}|$ (28)
$\leq\frac{1}{{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|,$ (29)
where $C_{i}={\frac{3}{c}}\cdot\frac{\sigma_{n}(r_{i})^{2}}{\gamma_{i}^{2}(1-2\alpha)}$ comes from Lemma 1 and we denote $\mathbf{C}=[C_{1},\ldots,C_{n}]$. To bound $\mathbf{R}(t)$, we decompose it using an intermediate time step $T=t-{\log(t)/c_{2}}>0$, where $c_{2}=1-\lambda_{2}(2)>0$. Let $T^{*}=\min\left\{t>1\,\middle|\,c_{2}t>\log(t)\right\}$. We obtain for all $t>T^{*}$,
@@
Using Lemma 1, we derive for $s>1$, $|\Delta{w}_{k}(s)|\leq\frac{2C_{k}}{{s-1}}$. We obtain $\|\mathbf{S}(1)\|\leq\|\mathbf{C}\odot\mathbf{X}\|$ and for $s>1$,
$\|\mathbf{S}(s)\|\leq\sqrt{\sum_{k=1}^{n}\frac{4C_{k}^{2}X_{k}^{2}}{(s-1)^{2}}}\leq\frac{2}{{s-1}}\|\mathbf{C}\odot\mathbf{X}\|.$
Applying this bound and using $c_{2}=1-\lambda_{2}(2)$ yields:
$\|\mathbf{R}(t)\|$ $\leq\lambda_{2}(2)^{t-T}\sum_{s=1}^{T}\lambda_{2}(2)^{T-s}\|\mathbf{S}(s)\|+\sum_{s=T+1}^{t}\lambda_{2}(2)^{t-s}\|\mathbf{S}(s)\|$
$\leq e^{-c_{2}(t-T)}\frac{2}{1-\lambda_{2}(2)}\|\mathbf{C}\odot\mathbf{X}\|+\frac{1}{T}\frac{2}{(1-\lambda_{2}(2))}\|\mathbf{C}\odot\mathbf{X}\|$
$\leq\left(\frac{2}{c_{2}t}+\frac{2}{{c_{2}t-\log(t)}}\right)\|\mathbf{C}\odot\mathbf{X}\|.$
@@
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\frac{1}{{t}}\left\|\mathbf{C}\odot\mathbf{X}\right\|+\frac{4}{{ct}}\|\mathbf{C}\odot\mathbf{X}\|+\frac{2}{\frac{c}{2}t-\log(t)}\|\mathbf{C}\odot\mathbf{X}\|,$
where $c=2c_{2}$. This bound simplifies to
$\left\|\mathbb{E}[\boldsymbol{Z}(t)]-\bar{x}_{\alpha}\mathbf{1}_{n}\right\|\leq\left(\frac{5}{{ct}}+\frac{4}{ct-2\log(t)}\right)\|\mathbf{C}\odot\mathbf{X}\|.$
Moreover, we have $\|\mathbf{C}\odot\mathbf{X}\|={\frac{{3}}{c(1-2\alpha)}}\|\mathbf{K}\odot\mathbf{X}\|$ where $\mathbf{K}=\left[\frac{\sigma_{n}(r_{k})^{2}}{\gamma_{k}^{2}}\right]_{k=1}^{n},$ which completes the proof. ∎